OpenStudy (anonymous):

At Central High School, there are 48 football players and 20 cheerleaders. What is the ratio of cheerleaders to football players? 12:5 1:48 5:12 48:1 @paki @op13456 @DragonFlame @NeedHelpQuick546 @mjoxprm360

OpenStudy (anonymous):

@scarymary101 @hartnn @shiraz14

OpenStudy (anonymous):

ok so there are 48 football players and 20 cheerleaders the ratio would be 48:28, just simplify down and you will get the answer

OpenStudy (anonymous):

shes right

OpenStudy (anonymous):

or excuse me the ratio is 20:48

OpenStudy (anonymous):

so its 5/12

OpenStudy (anonymous):

correct @yo_its__nate

OpenStudy (texaschic101):

cheerleaders: football players = 20/48 = 5/12

OpenStudy (anonymous):

horray

OpenStudy (texaschic101):

oops...5:12, put it in this form

OpenStudy (anonymous):

can someone give me a medal

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

anybody else want to give me a medal